at 20C a cell with a pressure potential of 3 bars is in equilibrium with the surrounding .4M solution of sucrose in an open beaker. what is the molar concentration of sucrose in the cell?
1
Expert's answer
2020-09-16T07:45:26-0400

 The solute potential "\Psi_S = -iCRT" , where i is the ionization constant, C is the molar concentration, R is the pressure constant (R = 0.0831 liter bars/mole K).

i = 1 (for sucrose)

"\Psi_{S(surround.)} = -(1)\times (0.4)\times (0.0831)\times (273+20)"

"\Psi_{S(surround.)} = -9.74"

"\Psi_{S(surround.)} = \Psi_P + \Psi_S"

"-9.74 = 3 + \Psi_S"

"\Psi_S = -12.74"

"\Psi_S = iCRT"

"-12.74 = -(1)\times C\times (0.0831)\times (273+20)"

"-12.74 = -24.34C"

"C = 0.52\; M"

Answer: 0.52 M